FREQUENCY COMB-REFERENCED MEASUREMENTS OF SELF- AND NITROGEN-PERTURBED LINE SHAPES IN THE ν_{1} + ν_{3} BAND OF ACETYLENE

\underline{MATTHEW J. CICH}, GARY V. LOPEZ, TREVOR J. SEARS^{a}, Department of Chemistry, Stony Brook University, Stony Brook, New York 11794; C. P. MCRAVEN, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973; A. W. MANTZ, Department of Physics, Astronomy, and Astrophysics, Connecticut College, New London, CT 06320; and DANIEL HURT-MANS, Service de Chimie Quantique et de Photophysique(Atoms, Molecules et Atmospheres), Universite Libre de Bruxelles, Bruxelles, Belgium B-10050.

Frequency comb technology has the potential to dramatically improve precision and accuracy in measured spectra, but few applications have yet been reported. One application that can benefit from the high stability of spectrometers referenced to frequency combs is the measurement of spectral line shapes. We have built an absorption spectrometer based on an extended cavity diode laser locked to an Er-fiber-based frequency comb operating near 1550 nm. Here we report the first measurements of line shapes using a frequency comb as a reference. We studied the P(11) line in the ν_{1} + ν_{3} combination band of acetylene at 195 739.649 513(8) GHz at several temperatures from 296K down to 175K. This talk focuses on the data taken at 296K. We used a hard collision model (Rautian-Sobel’man) fit to the measurements to determine selfand nitrogen- pressure broadening, pressure shift and Dicke narrowing parameters that are at least 2 orders of magnitude more precise than those reported in previous measurements. We compare these hard collision model results with those of two other widely-used models: the Galatry (soft collision) and the Voigt function. The temperature- and rotational leveldependence of these parameters have been measured and some of the results will be discussed in a separate talk. Acknowledgements:Acknowledgement is made to the Donors of the American Chemical Society Petroleum Research Fund for partial support of this research. CPM gratefully acknowledges support by DOE EPSCoR grant DOE-07ER46361 for work conducted at the University of Oklahoma. The measurements and analyses were performed under grants NNX09AJ93G and NNX08AO78G from the NASA Planetary and Atmospheres program.

aalso: Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973